# RiemannLoRA: A Unified Riemannian Framework for Ambiguity-Free LoRA Optimization

### v1

@@ Appendix A Inital point search
To derive (14), one may take $S$ to be block-diagonal with $S_{11}=\alpha I_{r},S_{22}=I_{r},\alpha\in\mathbb{R}$. ∎
Appendix B Dataset
B.1 Commonsense reasoning
@@
Following the approach of Hu et al. [2023], we combine the training datasets from all 8 tasks to form the final training set and evaluate performance on each task’s individual test dataset. We adopt queries structure from Hu et al. [2023] to Llama 3.2 instruct template. The prompt structure is demonstrated in Table 3.
@@
B.2 GSM8K and MetaMathQA
The model fine-tuning results on MetaMathQA dataset presented by Yu et al. [2024] are demonstrated in Section E.2. The full dataset can be found here: https://huggingface.co/datasets/meta-math/MetaMathQA. For fine-tuning, we used the MetaMathQA dataset, with validation conducted on GSM8K’s test set.
For training and validation, we employed the queries declared in Table 4.
Table 4: The structure of the queries for the MetaMathQA dataset
Task Role Fine-tuning Data Template
MetaMathQA system Below is an instruction that describes a math task.
Solve the problems and use "The answer is:" at the end of your answer.
user [QUERY]
assistant [RESPONSE]
GSM8K system Below is an instruction that describes a math task.
Solve the problems and use "The answer is:" at the end of your answer.
user [QUESTION]
assistant [ANSWER]
[…]
Appendix C Hyperparameters
For every method, we carefully preselected learning rate (optimization step size). For the RiemannLoRA with simulate_Adam flag enabled we also preselected parameters $\beta$ and $\gamma$ (see Algorithm 3). For the benchmark on the Commonsense reasoning dataset the hyperparameters are exhibited in Table 5 for SGD-like methods and in Table 6 for Adam-like methods. We selected $\beta=0.7,\gamma=0.9$ for the RiemannLoRA with simulate_Adam flag enabled without LOI and $\beta=0.85,\gamma=0.995$ for this algorithm with LOI.
@@ Appendix C Hyperparameters
Table 6: The parameters for different Adam variants (RiemannLoRA with simulate_Adam flag enabled) for fine-tuning on the Commonsense reasoning dataset
Optimizer learning rate
LoRA $2\cdot 10^{-4}$
LoRA-LOI $2\cdot 10^{-4}$
LoRA-QR $2\cdot 10^{-4}$
Pissa $5\cdot 10^{-5}$
RiemannLoRA $0.04$
RiemannLoRA-LOI $0.04$
For the benchmark on the MetaMathQA dataset the hyperparameters are exhibited in Table 7 for SGD-like and Adam-like methods simultaneously. We selected $\beta=0.7,\gamma=0.9$ for the RiemannLoRA with simulate_Adam flag enabled without LOI and $\beta=0.85,\gamma=0.995$ for this algorithm with LOI.
Table 7: The parameters for different SGD and Adam variants (RiemannLoRA with simulate_Adam flag disabled and enabled correspondingly) for fine-tuning on the MetaMathQA dataset
Optimizer learning rate
SGD Adam
LoRA $0.05$ $3\cdot 10^{-4}$
LoRA-LOI $0.08$ $7.5\cdot 10^{-5}$
DoRA $0.08$ $1\cdot 10^{-4}$
RiemannLoRA $0.08$ $0.02$
RiemannLoRA-LOI $0.08$ $0.01$
The non-tuned hyperparameters used for experiments on the Commonsense Reasoning and MetaMathQA datasets are presented in 8.
Table 8: Other non-tuned hyperparameter configurations for experiments on Commonsense reasoning and MetaMathQA datasets
Dataset Commonsense reasoning MetaMathQA
@@
Appendix D Subject-driven generation
To evaluate the effectiveness of our parameterization for subject-driven generation, two sets of experiments were conducted: one for the Adam optimizer and one for the SGD optimizer. Details for both sets of experiments are provided below.
@@ Appendix D Subject-driven generation
We used Stable Diffusion-2-base model for all experiments. For the Adam optimizer setup, we used a learning rate of 2e-5 for LoRA and 2e-3 for our method. The difference in learning rates is due to the fact that our method uses exponentially smoothed gradient norms as the second momentum, meanwhile the vanilla Adam uses elementwise matrix division by the second momentum. The betas = $(0.9,0.999)$ are identical for both methods. For the SGD setup, we used the same learning rate of 10 for both LoRA and our method. We set $q=15$ and $p$ = rank in all variations of our approach. We used a batch size of $4$ in all our experiments.
Figure 4: Comparison of text and image similarities for LoRA and our method at different training steps with SGD optimizers. Larger dots indicate a higher training step
@@
We used the DreamBooth dataset, which contains 25 different prompts and 30 various concepts. Due to computational costs, we only used half of the proposed concepts: can, candle, cat, cat2, colorful_sneaker, dog2, dog3, dog5, dog6, dog7, dog8, fancy_boot, grey_sloth_plushie, pink_sunglasses, vase. To measure the similarity between the original concept and the generated images, we used BaseIS: we synthesized 30 images for the base prompt <<a photo of a V*>> and measured the average pairwise cosine similarity with reference photos of the concept using the CLIP ViTB/32 model. To check the correspondence between the generated images and the textual prompts, we calculated Text Similarity (TS): we evaluated each concept with each of 25 prompts, synthesizing 10 images per prompt, and calculating the average pairwise cosine similarity with the prompts using the CLIP ViTB/32 model.
Figure 5: Additional visual comparison of our method and LoRA in Adam optimizers setup, checkpoint $600$
Figure 6: Additional visual comparison of our method and LoRA in Adam optimizers setup, checkpoint $1400$
Figure 7: Additional visual comparison of our method and LoRA in SGD optimizers setup, checkpoint $1000$
Results
Figures $2$ and $3$ (from the main text) represent the quantitative and qualitative results for the Adam setup. Figures 5 and 6 also present additional visual results for the Adam setup with checkpoints at $600$ and $1400$ steps. Figure 4 shows the metrics for the SGD setup experiments. In this figure we do not report metrics at the $2000$ step due to both original LoRA and our method starting to generate noisy images and dropping metrics significantly. Figure 7 contains a visual comparison of the SGD setup for various prompts and concepts at $1000$ step. Both the visual and numerical results show that our method outperforms LoRA in both the Adam and SGD setups. This is especially evident in experiments with SGD, which naturally generalizes to the Riemannian setting. However, there is also an increase in image similarity and faster concept learning in the Adam setup.
Appendix E Additional experimental results
E.1 Ablation study on initialization
The fine-tuning optimization was carried out by AdamW (Loshchilov and Hutter [2019]) First of the first, for every approach tested we preselected a suitable optimization step sizes. The list of all selected hyper-parameters for each method is specified in the Appendix C. Table 9 contains the accuracy of the trained model’s responses on the test dataset. The notation <<LoRA-A>> in the table means the utilization of the vanilla LoRA with non-zero $A$, the notation <<LoRA-B>> means the same with non-zero $B$, the notations <<Stiefel-A>> and <<Stiefel-B>> indicate vanilla LoRA with orthonormal initialization (for matrix $A$ and $B$ respectively). The naming of <<Stiefel-both>> involves taking the factors $A,B$ with orthonormal columns and with initialization like in (1).
Table 9: The average accuracy among $8$ tasks of fine-tuned Llama 3.2-1b via Adam using different LoRA initialization variants, tested on Commonsense reasoning benchmark. <<LOI>> stands for locally optimal initialization. LoRA rank is set to 16.

### v2

@@ Appendix A Inital point search
To derive the version that is utilized in practice, one may take $S$ to be block-diagonal with $S_{11}=\alpha I_{r},S_{22}=I_{r},\alpha\in\mathbb{R}$. ∎
Appendix B Riemannion with LOI
The complete framework is summarized in Algorithm 5, which consists of an LOI initialization step followed by max_iters iterations of the Riemannion optimizer.
Specifically, the first $3$ lines are dedicated to the computation of LOI via BackPropRSVD for given layer weights and oversampling parameter $p$, that increases the accuracy of the singular approximation. The $4$-th step initializes Heavy-Ball momentum matrices. Therefore, the asymptotical complexity of this stage is determined by the complexity of Algorithm 3.
Algorithm 5 Riemannion with LOI
Require: Weights $W\in\mathbb{R}^{m\times n}$, rank $r\in\mathbb{N}$, step size $\eta$, momentum coefficient $\beta$, weight decay coefficient $\gamma$, oversampling parameter $p$, power-step parameter $q$.
Ensure: Tuning parameters $\Delta W^{*}\in\mathcal{M}_{r}$.
Function:
[…]
5:for $i:=0,\dots,\mathtt{max\_iters}$ do
7:  $\dot{A}:=\nabla_{Z_{1}}\mathcal{L}\left(W^{\prime}+Z_{1}B_{R}^{\top}+A_{L}Z_{2}^{\top}\right)|_{Z_{1}=0,Z_{2}=B}.$ // $\nabla_{W}\mathcal{L}(W)B_{R}$
8:  $\dot{B}:=\nabla_{Z_{2}}\mathcal{L}\left(W^{\prime}+Z_{1}B_{R}^{\top}+A_{L}Z_{2}^{\top}\right)|_{Z_{1}=0,Z_{2}=B}.$ // $\nabla_{W}\mathcal{L}(W)^{\top}A_{L}$
9:  $\dot{A}_{\mathrm{prev}},\dot{B}_{\mathrm{prev}}:=\mathtt{ProjectLR}\left((A_{\mathrm{M}},B_{\mathrm{M}}),A_{L},B_{R}\right)$. // $\mathcal{O}\left((m+n)r^{2}\right)$
10:  $\dot{A},\dot{B}:=\beta\dot{A}_{\mathrm{prev}}+(I-A_{L}A_{L}^{\top})\dot{A},\;\beta\dot{B}_{\mathrm{prev}}+\dot{B}$ // $\mathcal{O}\left((n+m)r^{2}\right)$
11:  $\dot{A},\dot{B}:=\mathtt{ProjectLR}\left(\mathtt{OrthoLR}\left(A_{L},B_{R},\dot{A},\dot{B},r,\right),A_{L},B_{R}\right)$ // $\mathcal{O}\left((m+n)r^{2}+r^{3}\right)$
12:  $U,\Sigma,V^{\top}:=\mathtt{RetractionLR}\left(\left[-\eta\dot{A},A_{L}\right],\left[B_{R},-\eta(\dot{B}+\gamma B_{R})\right]\right)$ // $\mathcal{O}\left((m+n)r^{2}+r^{3}\right)$
13:  $A_{\mathrm{HB}},B_{\mathrm{HB}}:=[\dot{A},A_{L}],\;[B_{R},\dot{B}]$
14:  $A_{L},B:=U,\Sigma V^{\top}$ // $\mathcal{O}\left(nr^{2}\right)$
15:return $A_{L},B$.
Appendix C Riemann-SGD Algorithm
In Algorithm 6, we present the RiemannSGD version of the fine-tuning algorithm. The first steps are dedicated to the computation of the optimal initialization using BackPropRSVD(Algorithm 3). After initialization, the algorithm follows an Adam-like procedure adapted to the Riemannian setting: it maintains exponentially smoothed estimates of momentum and gradient norms ($4th$ step). The overall cycle encapsulates the computation of the Riemannian gradient at the current point(steps 787-8), the update of the adaptive momentum terms via ProjectLR(Algorithm 2), and the retraction step that projects the updated tangent direction back onto the manifold(steps 151715-17). This approach has asymptotical complexity of $\mathcal{O}((m+n)r^{2}+r^{3})$ and $(2(q+1)+\mathtt{max\_iters})$ amount of backward calls.
Algorithm 6 RiemannSGD with LOI
Ensure: Tuning parameters $\Delta W^{*}\in\mathcal{M}_{r}$.
Function:
[…]
5:for $i:=0,\dots,\mathtt{max\_iters}$ do
9:  $\dot{A}_{\mathrm{prev}},\dot{B}_{\mathrm{prev}}:=\mathtt{ProjectLR}\left((A_{\mathrm{HB}},B_{\mathrm{HB}}),A_{L},B_{R}\right).$ // $\mathcal{O}\left((m+n)r^{2}\right)$
11:  if $\mathtt{simulate\_Adam}$ then
14:  $U,\Sigma,V^{\top}:=\mathtt{RetractionLR}\left(\left[\eta\dot{A},A_{L}\right],\left[B_{R},\eta\dot{B}+B\right]\right)$ // $\mathcal{O}\left((m+n)r^{2}+r^{3}\right)$
15:  $A_{\mathrm{HB}},B_{\mathrm{HB}}:=[\dot{A},A_{L}],\;[B_{R},\dot{B}]$
16:  $A_{L},B:=U,\Sigma V^{\top}$ // $\mathcal{O}\left(nr^{2}\right)$
17:return $A_{L},B$.
Table 2: The average accuracy (in %) among 88 tasks of fine-tuned Llama 3.2-1b using different SGD variants, tested on Commonsense Reasoning benchmark. The <<R->> prefix stands for Riemannian, the postfix <<-LOI>> stands for locally optimal initialization. LoRA rank is set to 16. In all experiments excpet for <<-LOI>> the $A$ factor in $AB^{\top}$ has orthonormal columns and $B$ is zero.
Task BoolQ PIQA SIQA hella- wino- ARC-E ARC-C OBQA All
Initialization swag grande
Raw $40.1$ $55.4$ $50.3$ $25.8$ $50.0$ $61.9$ $41.8$ $42.8$ $46.0$
RSLoRA $64.5_{\pm 0.2}$ $77.2_{\pm 0.3}$ $68.1_{\pm 0.8}$ $86.3_{\pm 0.1}$ $58.2_{\pm 0.5}$ $76.5_{\pm 1.5}$ $58.6_{\pm 2.9}$ $70.0_{\pm 1.9}$ $70.0_{\pm 0.8}$
RiemannLoRA $\mathbf{65.1}_{\pm 0.4}$ $77.7_{\pm 0.3}$ $69.1_{\pm 1.5}$ $85.8_{\pm 0.2}$ $58.5_{\pm 0.4}$ $74.4_{\pm 1.9}$ $56.9_{\pm 1.3}$ $69.3_{\pm 2.5}$ $69.6_{\pm 0.9}$
RiemannLoRA-LOI $\mathbf{65.2}_{\pm 0.4}$ $\mathbf{79.4}_{\pm 0.4}$ $\mathbf{75.6}_{\pm 0.4}$ $\mathbf{87.3}_{\pm 0.1}$ $\mathbf{62.4}_{\pm 0.4}$ $\mathbf{79.9}_{\pm 0.8}$ $\mathbf{63.6}_{\pm 1.9}$ $\mathbf{73.8}_{\pm 0.5}$ $\mathbf{73.4}_{\pm 0.3}$
Appendix D LMO bounds
Let $A\in\mathcal{T}_{W}\mathcal{M}_{r}$ be a vector from the tangent space. Consider the problem
$\begin{split}&\langle A,X\rangle\to\max_{X},\\ &\text{s.t.}\quad\|X\|_{2}\leq 1,\end{split}$ (21)
whose (any) solution will be denoted by $X_{\ast}=\mathtt{Ortho}(A)$. Denote $f_{\ast}=\langle A,X_{\ast}\rangle$ the optimal value.
Consider the restricted (LMO) problem
$\begin{split}&\langle A,X\rangle\to\max_{X},\\ &\text{s.t.}\quad\|X\|_{2}\leq 1,\\ &X\in\mathcal{T}_{W}\mathcal{M}_{r}.\end{split}$ (22)
Let $X^{\ast}$ denote an optimizer of the task derived in (LABEL:eq:task2) and $f^{\ast}=\langle A,X^{\ast}\rangle$ its optimal value.
Proposition D.1.
Assume $A\in\mathcal{T}_{W}\mathcal{M}_{r}$ and let $P$ be the orthogonal projector onto $\mathcal{T}_{W}\mathcal{M}_{r}$. If
$\|X_{\ast}-PX_{\ast}\|_{2}\leq\varepsilon$
for some $\varepsilon\geq 0$, then
$\frac{1}{1+\varepsilon}\,f_{\ast}\leq f^{\ast}\leq f_{\ast}.$
Proof.
Since $A\in\mathcal{T}_{W}\mathcal{M}_{r}$, we have $PA=A$ and $P=P^{\top}$. By assumption $\|X_{\ast}-PX_{\ast}\|_{2}\leq\varepsilon$ and, because $\|X_{\ast}\|_{2}=1$, the reverse triangle inequality yields
$\big|\|X_{\ast}\|_{2}-\|PX_{\ast}\|_{2}\big|\leq\varepsilon,$
hence
$1-\varepsilon\leq\|PX_{\ast}\|_{2}\leq 1+\varepsilon.$
Define the normalized matrix
$Y:=\frac{PX_{\ast}}{\|PX_{\ast}\|_{2}}\in\mathcal{T}_{W}\mathcal{M}_{r},\qquad\|Y\|_{2}=1.$
By definition of $f^{\ast}$ (the maximum of $\langle A,\cdot\rangle$ over $\{X\in\mathcal{T}_{W}\mathcal{M}_{r}:\|X\|_{2}\leq 1\}$) we have
$f^{\ast}\geq\langle A,Y\rangle=\frac{\langle A,PX_{\ast}\rangle}{\|PX_{\ast}\|_{2}}=\frac{\langle PA,X_{\ast}\rangle}{\|PX_{\ast}\|_{2}}=\frac{\langle A,X_{\ast}\rangle}{\|PX_{\ast}\|_{2}}=\frac{f_{\ast}}{\|PX_{\ast}\|_{2}}.$
Using the inequality $\|PX_{\ast}\|_{2}\leq 1+\varepsilon$ we obtain the claimed lower bound
$f^{\ast}\geq\frac{f_{\ast}}{1+\varepsilon}.$
The upper bound $f^{\ast}\leq f_{\ast}$ is trivial because the feasible set of the task derived in (LABEL:eq:task2) is a subset of the feasible set of the problem described in (LABEL:eq:task1). Combining the two inequalities gives
$\frac{1}{1+\varepsilon}\,f_{\ast}\leq f^{\ast}\leq f_{\ast},$
which completes the proof. ∎
Proposition D.2.
(Upper bound on the projection residual.) With the notation above, the residual
$\varepsilon:=\|X_{\ast}-PX_{\ast}\|_{2}$
satisfies $\varepsilon\leq 1$.
Proof.
Write the thin SVD of the base point $W$ as $W=U_{r}\Sigma_{r}V_{r}^{\top}$, where $U_{r}\in\mathtt{St}(m,r)$ and $V_{r}\in\mathtt{St}(n,r)$ have $r$ orthonormal columns. Let $U_{\perp}$ and $V_{\perp}$ be orthonormal complements so that $U_{r}\;U_{\perp}]\in\mathbb{R}^{m\times 2r}$ and $V_{r}\;V_{\perp}]\in\mathbb{R}^{n\times 2r}$ are orthogonal matrices (here we assume $m,n\geq 2r$; if $m$ or $n$ is smaller the argument is adapted in the obvious way).
Because $A\in\mathcal{T}_{W}\mathcal{M}_{r}$, one can write (using a suitable $2r\times 2r$ block representation) (Vandereycken, 2013)
$A=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\begin{bmatrix}\tilde{\mathcal{A}}&\tilde{\mathcal{B}}\\[4.0pt] \tilde{\mathcal{C}}&0\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\[2.0pt] V_{\perp}^{\top}\end{bmatrix}$,
where $\tilde{\mathcal{A}},\tilde{\mathcal{B}},\tilde{\mathcal{C}}\in\mathbb{R}^{r\times r}$ (the $(2,2)$ block vanishes for tangent-space elements of this form). The rank of $A$ is therefore at most $2r$.
Let $X_{\ast}=\mathtt{Ortho}(A)$ be a maximizer of (LABEL:eq:task1); represent $X_{\ast}$ in the same block basis,
$X_{\ast}=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\begin{bmatrix}\mathcal{A}&\mathcal{B}\\[4.0pt] \mathcal{C}&\mathcal{D}\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\[2.0pt] V_{\perp}^{\top}\end{bmatrix}$,
where the $2r\times 2r$ block matrix is orthogonal,
$\begin{bmatrix}\mathcal{A}&\mathcal{B}\\[4.0pt] \mathcal{C}&\mathcal{D}\end{bmatrix}\in\mathcal{O}(2r),$ (23)
where $\mathcal{O}(2r)$ denotes the orthogonal group in $\mathbb{R}^{2r\times 2r}$. By Lemma D.1 below we have
$PX_{\ast}=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\begin{bmatrix}\mathcal{A}&\mathcal{B}\\[4.0pt] \mathcal{C}&0\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\[2.0pt] V_{\perp}^{\top}\end{bmatrix}$.
Therefore
$X_{\ast}-PX_{\ast}=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\begin{bmatrix}0&0\\[4.0pt] 0&\mathcal{D}\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\[2.0pt] V_{\perp}^{\top}\end{bmatrix}=U_{\perp}\mathcal{D}V_{\perp}^{\top},$
and hence
$\|X_{\ast}-PX_{\ast}\|_{2}=\|\mathcal{D}\|_{2}.$
Because the block matrix in (23) is orthogonal, its lower block
$\begin{bmatrix}\mathcal{C}\\[2.0pt] \mathcal{D}\end{bmatrix}\in\mathbb{R}^{2r\times r}$
has orthonormal columns. Thus $\big\|\begin{smallmatrix}\mathcal{C}\\[2.0pt] \mathcal{D}\end{smallmatrix}\big\|_{2}=1$, and in particular $\|\mathcal{D}\|_{2}\leq 1$. Consequently $\varepsilon=\|X_{\ast}-PX_{\ast}\|_{2}\leq 1$, as claimed. ∎
Lemma D.1.
Let
$W=U_{r}\Sigma_{r}V_{r}^{\top},\quad U_{r}\in\mathtt{St}(m,r),\;V_{r}\in\mathtt{St}(n,r),$
and write an arbitrary matrix $X\in\mathbb{R}^{m\times n}$ in the block basis determined by $U_{r}\;U_{\perp}]$ and $V_{r}\;V_{\perp}]$,
$X=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\begin{bmatrix}\mathcal{A}&\mathcal{B}\\[4.0pt] \mathcal{C}&\mathcal{D}\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\[2.0pt] V_{\perp}^{\top}\end{bmatrix}$.
Then the orthogonal projector $P$ onto $\mathcal{T}_{W}\mathcal{M}_{r}$ acts as
$PX=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\begin{bmatrix}\mathcal{A}&\mathcal{B}\\[4.0pt] \mathcal{C}&0\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\[2.0pt] V_{\perp}^{\top}\end{bmatrix}$.
Proof.
One convenient expression for the orthogonal projector onto the tangent space at $W=U_{r}\Sigma_{r}V_{r}^{\top}$ is
$P(\cdot)=(\cdot)-(I-U_{r}U_{r}^{\top})\,(\cdot)\,(I-V_{r}V_{r}^{\top}).$
Applying this to $X$ written in the chosen block basis yields
$PX$ $=X-(I-U_{r}U_{r}^{\top})X(I-V_{r}V_{r}^{\top})=$
$=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\begin{bmatrix}\mathcal{A}&\mathcal{B}\\[4.0pt] \mathcal{C}&\mathcal{D}\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\[2.0pt] V_{\perp}^{\top}\end{bmatrix}-\begin{bmatrix}0&U_{\perp}\end{bmatrix}\begin{bmatrix}\mathcal{A}&\mathcal{B}\\[4.0pt] \mathcal{C}&\mathcal{D}\end{bmatrix}\begin{bmatrix}0\\[2.0pt] V_{\perp}^{\top}\end{bmatrix}$
$=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\left(\begin{bmatrix}\mathcal{A}&\mathcal{B}\\[4.0pt] \mathcal{C}&\mathcal{D}\end{bmatrix}-\begin{bmatrix}0&0\\[4.0pt] 0&\mathcal{D}\end{bmatrix}\right)\begin{bmatrix}V_{r}^{\top}\\[2.0pt] V_{\perp}^{\top}\end{bmatrix}$
$=\begin{bmatrix}U_{r}&U_{\perp}\end{bmatrix}\begin{bmatrix}\mathcal{A}&\mathcal{B}\\[4.0pt] \mathcal{C}&0\end{bmatrix}\begin{bmatrix}V_{r}^{\top}\\[2.0pt] V_{\perp}^{\top}\end{bmatrix}$,
which proves the lemma. ∎
Remark D.1.
In our experiments we observed that the residual $\varepsilon$ is typically small (in our reported runs it did not exceed $0.1$).
Appendix E Dataset
E.1 Commonsense reasoning
@@
Following the approach of Hu et al. (2023), we combine the training datasets from all 8 tasks to form the final training set and evaluate performance on each task’s individual test dataset. We adopt queries structure from Hu et al. (2023) to Llama 3.2 instruct template. The prompt structure is demonstrated in Table 3.
@@
Appendix F Hyperparameters
For each optimization method, we carefully preselected the learning rate (step size). For the Riemannion method, we set the momentum to $0.9$, and for the Muon method, we set the momentum to $0.95$. In both cases, we additionally tuned the weight-decay hyperparameter (see Algorithm 4). The final choices of these hyperparameters are summarized in Table 4.
In the Commonsense reasoning benchmark, the hyperparameters are reported in Table 5 for the SGD-like methods and in Table 4 for the Adam-like methods. For the LOI method, we selected the following hyperparameters: the backprop RSVD oversampling parameter was set to $r_{k}=16$, the backprop powerstep parameter was set to $q=1$, and the multiplier $\alpha$ was set to $\frac{-0.01}{\sqrt{r}}$.
Table 4: The parameters for different Adam variants for fine-tuning on the Commonsense reasoning dataset
Optimizer learning rate weight decay
Adam $0.0002$ 0.01
DoRA $0.0003$ 0.01
Muon $0.0005$ 0.2
DoneRITE $0.0005$ 0.0001
RPrecAdamW $0.0005$ 0.5
Riemannion $0.0001$ 0.00316
@@ Appendix F Hyperparameters
The non-tuned hyperparameters used for experiments on the Commonsense Reasoning dataset are presented in 6.
Table 6: Other non-tuned hyperparameter configurations for experiments on Commonsense reasoning dataset
Dataset Commonsense reasoning
@@
Appendix G Subject-driven generation
@@
We used Stable Diffusion-2-base model with a batch size of $4$ for all experiments. For both methods, we set the betas to $0.9$ and $0.999$ and the weight decay to $0.1$. In all variations of our approach, we set $q=15$, $p$ = rank and
$\alpha=-\frac{1}{\sqrt{\text{rank}}}$. We used a learning rate of 2e-5 to train both our and the LoRA models, which were used to generate images shown in Figures 2 and 3.
@@
We used the DreamBooth dataset, which contains 25 different prompts and 30 various concepts. Due to computational costs, we only used half of the proposed concepts to evaluate metrics: can, candle, cat, cat2, colorful_sneaker, dog2, dog3, dog5, dog6, dog7, dog8, fancy_boot, grey_sloth_plushie, pink_sunglasses, vase. To measure the similarity between the original concept and the generated images, we used Image Similarity (IS): we synthesized 30 images for the base prompt <<a photo of a V*>> and 10 images for each of 25 editing prompts (like <<a V* on top of a dirt road>>). We then measured the average pairwise cosine similarity with reference photos of the concept using the DINO model. To check the correspondence between the generated images and the textual prompts, we calculated Text Similarity (TS): we evaluated each concept with each of 25 prompts, synthesizing 10 images per prompt, and calculating the average pairwise cosine similarity with the prompts using the CLIP ViTB/32 model.
Figure 3: Additional visual comparison of our method and LoRA, checkpoint $600$
Additional results
Figure 3 shows an additional visual comparison of our method and LoRA after 600 training steps. As can be seen, our method learns the concept much faster than the original LoRA, while preserving editing capabilities.
Appendix H Additional experimental results
H.1 Ablation study on initialization
The fine-tuning optimization was carried out by AdamW (Loshchilov & Hutter (2019)) First of the first, for every approach tested we preselected a suitable optimization step sizes. The list of all selected hyper-parameters for each method is specified in the Appendix F. Table 7 contains the accuracy of the trained model’s responses on the test dataset. The notation <<LoRA-A>> in the table means the utilization of the vanilla LoRA with non-zero $A$, the notation <<LoRA-B>> means the same with non-zero $B$, the notations <<Stiefel-A>> and <<Stiefel-B>> indicate vanilla LoRA with orthonormal initialization (for matrix $A$ and $B$ respectively). The naming of <<Stiefel-both>> involves taking the factors $A,B$ with orthonormal columns and with initialization like in (15).
Table 7: The average accuracy among 88 tasks of fine-tuned Llama 3.2-1b via Adam using different LoRA initialization variants, tested on Commonsense reasoning benchmark. <<LOI>> stands for locally optimal initialization. LoRA rank is set to 16.